you have 12 mp3 songs on your computer. 1/4 of them don't work. How many mp3's do you have to buy to replace the ones that don't work.
Answer by nerdybill(7384)

(1/4) 12 = 12/4 = 3
This means 3 mp3 songs don't work.
.
Therefore, you'll need to buy 3 mp3 songs to replace them.